OpenStudy (anonymous):

Jamie and Nicole are buying streamers for their mother’s surprise birthday party. They need to buy purple and gold streamers. The purple streamers cost $2 per roll and the gold streamers cost $3 per roll. They have $25 to spend on streamers. Which of the following is a possible combination of streamers? A. 1 roll of purple streamer and 8 rolls of gold streamer B. 3 rolls of purple streamer and 6 rolls of gold streamer C. 7 rolls of purple streamer and 4 rolls of gold streamer D. 9 rolls of purple streamer and 3 rolls of gold streamer

OpenStudy (texaschic101):

2p + 3g = 25 now plug in your answer choices and see which one works

OpenStudy (texaschic101):

I will do the first one for you... 2p + 3g = 25 when p = 1 and g = 8 now sub 2(1) + 3(8) = 25 2 + 24 = 25 26 = 25 (incorrect..this one will not work)

OpenStudy (anonymous):

hmmm let me try C

OpenStudy (anonymous):

C=29

OpenStudy (texaschic101):

C won't work...try B

OpenStudy (anonymous):

b= 21

OpenStudy (texaschic101):

that works..can there be more then 1 answer ?

OpenStudy (anonymous):

No

OpenStudy (texaschic101):

b = 2(3) + 3(6) = 6 + 18 = 24 <-- correct answer

OpenStudy (texaschic101):

answer is B

OpenStudy (anonymous):

thanks

OpenStudy (texaschic101):

anytime :)